please help me solve this problem in algrebra --number of elements for two sets
in a school, 37% of the pupils study principles of accounts, and 28% study principles of Business. If 15% study both business subjects, what percentage of the pupils:
a) do not study either
b) study principles of accounts only
c) study principles of business only
email address: bccastlegirl@live.com

Answer by stanbon(75887)

in a school, 37% of the pupils study principles of accounts, and 28% study principles of Business. If 15% study both business subjects, what percentage of the pupils:
Draw the Venn Diagram of the data:
Put 15 in the intersection.
Put 37-15 = 22 in the remaining part of prin of accounts
Put 28-15 = 13 in the remaining part of prin of business
a) #(do not study either) = 100-(22+15+13) = 50%
-------
b) #(study principles of accounts only)= 22%
-------
c) #(study principles of business only) = 13%
=================================================
Cheers,
Stan H.
================